# Phase transition near the Big Bang in the lattice theory of gravity and some cosmological considerations

### v6

@@ III Superhigh temperatures in the lattice model of gravity
Now let’s simulate the superhigh temperature in the lattice theory. As is known, in quantum theory, the partition function $\mathop{\rm tr}\nolimits\mathop{\rm exp}\nolimits(-\beta{\cal H})$ differs from the transition amplitude over a finite time $\Delta t$, represented as a functional integral, by the transition to imaginary time $\Delta t=-i\beta$, $\beta=1/T$, and taking the trace. We must study some properties of the partition function (10) in the case of ultra-high temperatures, that is, $\beta\longrightarrow 0$.
@@
Suppose that a D-dimensional lattice has two (D-1)-dimensional sublattices $\Sigma_{1}$ and $\Sigma_{2}$ that form its boundary. For simplicity, we assume that between $\Sigma_{1}$ and $\Sigma_{2}$ there are $N<\infty$ 4D lattice layers, each one edge thick. Between $\Sigma_{1}$ and $\Sigma_{2}$ there is a 3D sublattice $\Sigma_{0}$, and the entire 4D lattice is symmetrical with respect to the mirror reflection relative to the $\Sigma_{0}$ sublattice. It follows that the sublattices $\Sigma_{1}$ and $\Sigma_{2}$ are identical, which makes it possible to calculate the partition function.
@@
Statement. In a certain finite neighborhood of the point $\beta=0$, the free energy of the partition function (10), with the exception of the term of the form $(-(C_{1}\cdot\mathop{{\mathfrak{M}}}\nolimits+4\mathop{{\mathfrak{N}}}\nolimits\nu_{\Psi})\cdot\ln\beta)$ (see (46)), is a holomorphic function of the variable $\beta$. All action symmetries (4), including discrete PT symmetries, are preserved. $\square$
Let us give some arguments in favor of the Statement. Let us consider the high-temperature expansion of the partition function in the 2D Ising model, which is the sum over all closed paths with self-intersections on the lattice. Let the lattice contain $\mathop{{\mathfrak{L}}}\nolimits\longrightarrow\infty$ edges and let a path (possibly several non-intersecting paths) with a total of $\mathop{{\mathfrak{l}}}\nolimits$ edges be fixed on it. Note that $\mathop{{\mathfrak{l}}}\nolimits$ is an even number and $\mathop{{\mathfrak{l}}}\nolimits\geq 4$. When calculating the partition function, it is convenient to isolate the factor $(\cosh\beta)^{\mathop{{\mathfrak{L}}}\nolimits}$, which does not give a singularity in the free energy. Then each edge of the path should be assigned a weight $(\tanh\beta)$. Let $g_{\mathop{{\mathfrak{l}}}\nolimits}$ be the number of such closed paths. We have a strict inequality
@@
This implies the validity of the Statement. In turn, it follows from the Statement that in the high-temperature phase no symmetries are broken, including discrete PT symmetry.
Based on the validity of the Statement, we will begin to calculate the averages of some operators.
@@
We have a chain of equalities:
@@
Here $N$ is a normalizing constant, the integral is calculated only over the tetrad and Dirac field variables, but not over the connection variables. If the integration also included an integral over the variables $\{\Omega\}$, then any gauge non-invariant quantity, like ${\Theta}^{a}_{{{\cal V}}_{1}{{\cal V}}_{2}}$, would vanish identically. But the integral (48) is meaningful, since the gauge is fixed in it. In (48), the first equality is the definition of the average value, the second equality is a consequence of the Statement, the third and fourth equalities are consequences of the equalities (47) and (16), respectively. The last equality follows from the mirror symmetry of the complex with respect to the 3D subcomplex $\Sigma_{0}$.
@@
The calculation of the vacuum mean value ${\Theta}^{a}_{{{\cal V}}_{1}{{\cal V}}_{2}}$ in the continuum limit in the Minkowski signature at zero temperature was carried out in the work Vergeles (2021b):
@@
Comparison of Eqs (49) and (50) shows that in the lattice theory of gravity coupled with Dirac field there is a phase transition in temperature.
Phase transition from the zero value of the bilinear form (with respect to Dirac fields) $\Theta^{a}_{\mu}$ to its nonzero value and the meaning of this transition for the 4D theory of gravity are discussed in Volovik (2022). In this regard, an article Volovik (1990) may also be of interest.
@@ IV Einstein equation and solution
is used, $H={\dot{a}}/a$ is Hubble constant and $a(t)$ is the cosmic scale factor. We assume $c=1$. Appropriate remarks are made in formulas with restored dimension.
@@ IV.1 Solution of Einstein’s equations with bare cosmological constant
Another equation $\nabla_{a}T^{a}_{b}=0$ is a consequence of equations (59), and therefore it does not need to be considered. With the help of the Hubble constant $H(t)$ Eqs. (59) are rewritten as follows:
@@
So we have 3 unknown functions $\{\tilde{\varepsilon}(t),\tilde{p}(t),H(t)\}$ and 2 equations (60). The missing equation is the equation of state relating energy density and pressure. Regarding the equation of state, the following facts are reliably known: (i) in the case of an ideal Boltzmann gas, we have $\tilde{p}=(2/3)\tilde{\varepsilon}$; (ii) in the case of real dusty matter, we have $\tilde{p}=0$; (iii) in the case of real ultrarelativistic matter we have $\tilde{p}=\tilde{\varepsilon}/3$; (iiii) in the case of vacuum energy and pressure, we have $\tilde{p}=-\tilde{\varepsilon}$. In all four cases, the energy density and pressure are linearly related. In addition, these quantities have the same dimension. Therefore, we propose to accept the following hypothesis:
@@
This equation is linear and inhomogeneous with an unknown function $\varkappa(t)$, the asymptotics of which are further determined based on the known dynamics. The set of equations (60) and (61) has a solution:
@@
Here $H_{0}$ is the integration constant which plays the role of the Hubble constant at the beginning of the inflation phase. The moment $t_{0}>0$ is the conditional moment of the beginning of the inflation phase.
@@
In (65) dimension is restored. Then $H_{0}t_{\mbox{inf}}\cong 100$. Let’s take $\varkappa(t_{0})\cong 1/150$. Assume that during the time $t_{\mbox{inf}}$ the function $\varkappa$ changes insignificantly. Then for $t_{0}<t<t_{0}+t_{\mbox{inf}}$ the solutions (62), (63), (64) take the form
@@ V Thermodynamic considerations
The purpose of this consideration is to (at least superficially) explain the state equation (61).
@@
acquires the meaning of local vacuum temperature. Here the dimension is restored. An exact calculation of the local vacuum temperature in de Sitter spacetime has been done in Volovik (2009, 2024, 2023): $T_{\mbox{vac}}=\hbar H/\pi$. Local temperatures differ from the horizon temperature in de Sitter space-time Gibbons and Hawking (1977), although all the temperatures discussed here are of the same order. As can be seen, simple qualitative reasoning within the framework of quantum field theory leads to a correct estimate of the local vacuum temperature.
@@
Let us determine the temperature value in Kelvin degrees at the beginning of the inflation process, when, according to some estimates $H_{0}\sim 10^{39}\mbox{s}^{-1}$. Then
$T_{0}\sim\frac{\hbar H_{0}}{k}\sim\left(10^{28}\right)^{\circ}\mbox{K}.$ (75)
Here $k$ is the Boltzmann constant. The temperature estimate (75) is within the known temperature estimates in the initial phase of inflation.
@@
The last relation is contained in Volovik (2023). According to (74) the vacuum temperature $T\approx\hbar H$. Therefore, the ratio (77) is rewritten as follows:
Statement. If we know the energy density as a function of temperature, then the free energy density is
Checking the Statement is reduced to checking the fact that the function (79) satisfies the following system of equations:
@@ VI Conclussion
was proposed for lattice theory of gravity coupled with Dirac field. The relation (82) is justified by the fact that the values under the average transform identically under all symmetry transformations. The relation (50), obtained by direct calculation in the continuum limit at zero temperature, is another argument in favor of the validity of the relation (82).
@@
The last equality means that in the lattice theory of gravity in the high-temperature phase, space collapses into a point.
Obviously, in the long-wavelength limit at zero temperature, the contribution of the Dirac field to the energy-momentum tensor is nonzero (50).
The above means that in the lattice theory of gravity there is a phase transition in temperature (possibly more than one). In the high-temperature phase, space collapses into a point, the mean of the energy-momentum tensor is zero and the PT symmetry is not broken. This statement is proved for 4D lattice model of gravity coupled with Dirac field. In the low-temperature phase, these quantities are nonzero and the PT symmetry is broken. The role of the order parameter is played by the mean $\langle e^{a}_{\mu}\rangle$ which becomes nonzero in the low-temperature phase. The process of exponential expansion of space begins, turning into a power law of expansion. During the phase transition from the high-temperature to the low-temperature phase, domains with opposite mean values $\langle e^{a}_{\mu}\rangle$ and $\langle\Theta^{a}_{\mu}\rangle$ can be formed. The domain wall separating such domains was studied in Vergeles (2021b)

### v9

@@ III Superhigh temperatures in the lattice model of gravity
Let us consider the state of lattice theory at ultrahigh temperatures. As is known, in quantum theory, the partition function $\mathop{\rm tr}\nolimits\mathop{\rm exp}\nolimits(-\beta{\cal H})$ differs from the transition amplitude over a finite time $\Delta t$, represented as a functional integral, by the transition to imaginary time $\Delta t=-i\beta$, $\beta=1/T$, and taking the trace. We must study some properties of the partition function (10) in the case of ultra-high temperatures, that is, $\beta\longrightarrow 0$.
@@
Suppose that a 4D lattice has two 3D sublattices $\Sigma_{1}$ and $\Sigma_{2}$ that form its boundary. For simplicity, we assume that between $\Sigma_{1}$ and $\Sigma_{2}$ there are $N<\infty$ 4D lattice layers, each one edge thick. The sublattices $\Sigma_{1}$ and $\Sigma_{2}$ are considered identical, that is, there is a one-to-one correspondence between all elements of these sublattices. The last property makes it possible to calculate the partition function.
@@
Statement 1. In a certain finite neighborhood of the point $\beta=0$, the free energy of the partition function (10), with the exception of the term of the form $(-(C_{1}\cdot\mathop{{\mathfrak{M}}}\nolimits+4\mathop{{\mathfrak{N}}}\nolimits\nu_{\Psi})\cdot\ln\beta)$ (see (42)), is a holomorphic function of the variable $\beta$. All symmetries of the action (4), including discrete PT symmetries, are preserved. $\square$
Let us give some arguments in favor of the Statement 1. Let us consider the high-temperature expansion of the partition function in the 2D Ising model, which is the sum over all closed paths with self-intersections on the lattice. Let the lattice contain $\mathop{{\mathfrak{L}}}\nolimits\longrightarrow\infty$ edges and let a path (possibly several non-intersecting paths) with a total of $\mathop{{\mathfrak{l}}}\nolimits$ edges be fixed on it. Note that $\mathop{{\mathfrak{l}}}\nolimits$ is an even number and $\mathop{{\mathfrak{l}}}\nolimits\geq 4$. When calculating the partition function, it is convenient to isolate the factor $(\cosh\beta)^{\mathop{{\mathfrak{L}}}\nolimits}$, which does not give a singularity in the free energy. Then each edge of the path should be assigned a weight $(\tanh\beta)$. Let $g_{\mathop{{\mathfrak{l}}}\nolimits}$ be the number of such closed paths. We have a strict inequality
@@
This implies the validity of the Statement 1. In turn, it follows from the Statement1 that in the high-temperature phase no symmetries are broken, including discrete PT symmetry.
Based on the validity of the Statement 1, we will begin to calculate the averages of some operators.
@@
According to the general rule, the integral does not change when the integration variables are changed. In our case, the integral over the fermionic and tetrad variables and the sum over the fermionic states $\xi$ are of interest. The integral over the connection variables $\{\Omega\}$ should be excluded. Otherwise, further reasoning will be meaningless. Let’s consider two integrals.
First integral:
$Z\{\Omega\}=\sum_{\xi}\int_{e}\int_{\Psi^{{\dagger}}\Psi}\Phi_{2\xi}^{{\dagger}}\mathop{\rm exp}\nolimits(\beta\mathop{{\mathfrak{A}}}\nolimits)\Phi_{1\xi}\stackrel{{\scriptstyle PT}}{{\longrightarrow}}Z\{\Omega\}.$
Second integral:
$\langle{\Theta}^{a}_{{{\cal V}}_{1}{{\cal V}}_{2}}\rangle_{e,\Psi}=N\sum_{\xi}\int_{e}\int_{\Psi^{{\dagger}}\Psi}\Phi_{2\xi}^{{\dagger}}{\Theta}^{a}_{{{\cal V}}_{1}{{\cal V}}_{2}}\mathop{\rm exp}\nolimits(\beta\mathop{{\mathfrak{A}}}\nolimits)\Phi_{1\xi}\quad\stackrel{{\scriptstyle PT}}{{\longrightarrow}}\quad-\langle{\Theta}^{a}_{{{\cal V}}_{1}{{\cal V}}_{2}}\rangle_{e,\Psi}.$
Let’s consider the second integral in more detail. We have a chain of equalities:
@@
Here $N\{\Omega\}$ is a normalizing constant, the integral is calculated only over the tetrad and Dirac field variables, but not over the connection variables. If the integration also included an integral over the variables $\{\Omega\}$, then any gauge-non-invariant quantity under the integral, like ${\Theta}^{a}_{{{\cal V}}_{1}{{\cal V}}_{2}}$, would turn into zero identically. But the integral (44) is meaningful, since the gauge is fixed in it. In (44), the first equality is the definition of the average value, the second equality is a consequence of the Statement 1, the third and fourth equalities are consequences of the equalities (43) and (16), respectively.
@@
The vacuum mean of ${\Theta}^{a}_{{{\cal V}}_{1}{{\cal V}}_{2}}$ in the long-wave limit in the Minkowski signature at zero temperature was calculated in Vergeles (2021b):
@@
A comparison of the equations (45) and (46) shows that in the lattice theory of gravity coupled with the Dirac field, there is a phase transition in temperature.
The phase transition from a phase with a zero value of the average value $\Theta^{a}_{\mu}$ to a phase with a non-zero value of this value and the physical significance of this transition were considered in the work Volovik (2022). The work of Volovik (1990) may also be interesting in this regard.
@@ IV Einstein equation and solution
is used, $H={\dot{a}}/a$ is Hubble parameter and $a(t)$ is the cosmic scale factor. We assume $c=1$. Appropriate remarks are made in formulas with restored dimension.
@@ IV.1 Solution of Einstein’s equations with bare cosmological constant
Another equation $\nabla_{a}T^{a}_{b}=0$ is a consequence of equations (55), and therefore it does not need to be considered. With the help of the Hubble parameter $H(t)$ Eqs. (55) are rewritten as follows:
@@
So we have 3 unknown functions $\{\tilde{\varepsilon}(t),\tilde{p}(t),H(t)\}$ and 2 equations (56). The missing equation is the equation of state relating energy density and pressure. The following facts are known regarding the equation of state of relativistic matter: (i) in the case of real dust-like matter $p=0$; (ii) in the case of real ultrarelativistic matter $p=\varepsilon/3$; (iii) in the case of the energy density and vacuum pressure in de Sitter space $p=-\varepsilon$. In all three cases, the energy density and pressure are linearly related. In addition, these quantities have the same dimensionality. Therefore, for the energy density and pressure, including vacuum values, we accept the following hypothesis of a linear relation:
@@
This equation is linear and inhomogeneous with an unknown dimensionless function $\varkappa(t)$, the asymptotics of which are further determined based on the known dynamics. The set of equations (56) and (57) has a solution:
@@
Here $H_{0}$ is the integration constant which plays the role of the Hubble parameter at the beginning of the inflation phase. The moment $t_{0}>0$ is the conditional moment of the beginning of the inflation phase.
From solutions (58)-(60) it is clear that case (i) corresponds to the value $\varkappa(t)=1$, case (ii) — the value $\varkappa(t)=4/3$, and case (iii) — the value $\varkappa(t)=0$.
@@
In (61) dimension is restored. Then $H_{0}t_{\mbox{inf}}\cong 100$.
Estimates of the Hubble parameter in the inflation phase differ greatly in different sources. We adopted the value $H\sim 1.7\cdot 10^{15}\mbox{GeV}$, which is equivalent to the value (61) Linde (1996). In the paper Starobinsky (2000) the value of the Hubble parameter at the end of the inflation phase is estimated as $\sqrt{G}H<10^{-5}$. This value corresponds to $H\sim 10^{38}\mbox{cm}^{-1}$, which is close to the value (61).
Let us take $\varkappa(t_{0})\cong 1/150$ and assume that during time $t_{\mbox{inf}}$ the function $\varkappa$ changes insignificantly. This assumption means (see the first of Eq. (58)) that in the inflation phase $|\dot{H}|\ll H^{2}$. The last inequality is a necessary condition for inflation Motohashi et al. (2015). Then in the time interval $t_{0}<t<t_{0}+t_{\mbox{inf}}$ the solutions (58), (59), (60) take the form
@@
Let us estimate the temperature $T_{c}$ of the phase transition that violates PT symmetry. For this, we use the result of Volovik Volovik (2009, 2024a, 2024b, 2023), who calculated the local temperature of the ”vacuum” in de Sitter space:
$T_{\mbox{vac}}=\frac{\hbar H}{\pi}.$ (65)
Although in the theory under study the Hubble parameter, unlike in the case of de Sitter space, is not constant, here we will adopt the formula (65) to estimate the temperature (see also Section V).
It has been shown that at the phase transition point the classical Einstein equations (56) are not valid. However, we use them only for a qualitative estimation. Since at the phase transition point the mean of the fermion energy-momentum tensor is zero ($\varepsilon=0$), then according to the second equation (56) we have $H_{c}\sim\sqrt{\Lambda_{0}}$. From here and using (65), (53) we obtain the estimate:
$T_{c}\sim\frac{\hbar c}{l_{P}}\sim 10^{18}\mbox{GeV},\quad\mbox{or}\quad T_{c}\sim 10^{31}\mbox{K}.$ (66)
The phase transition temperature can also be estimated as the energy of the Dirac sea enclosed in the Planck volume $V_{P}\sim l_{P}^{3}$: $T_{c}\sim(\hbar c/l_{P}^{4})l_{P}^{3}\sim\hbar c/l_{P}$. This temperature is of the same order of magnitude as the temperature of the Grand Unification.
Let us also estimate the temperature in degrees Kelvin at the beginning of the inflation phase, when according to (61) $H_{0}\sim 10^{39}\mbox{s}^{-1}$. Then
$T_{0}\sim\frac{\hbar H_{0}}{k}\sim 10^{28}\mbox{K}.$ (67)
@@ V Thermodynamic considerations
The purpose of this Section is to qualitatively understand the result of temperature calculation (65). It will be shown that the local vacuum temperature in a dynamic Universe with a changing scale factor is always of the order of $\hbar H$.
@@
acquires the meaning of local vacuum temperature. Here the dimension is restored. An exact calculation of the local vacuum temperature in de Sitter spacetime has been done in Volovik (2009, 2024a, 2024b, 2023): $T_{\mbox{vac}}=\hbar H/\pi$. From the method of calculating the temperature in the cited works it is evident that the calculation result is valid when $|\dot{H}|\ll H^{2}$. The last restriction takes place in the inflation phase. Local temperatures differ from the horizon temperature in de Sitter space-time Gibbons and Hawking (1977), although all the temperatures discussed here are of the same order. As can be seen, simple qualitative reasoning within the framework of quantum field theory leads to a correct estimate of the local vacuum temperature.
@@
The last relation is contained in Volovik (2023). It is the quantity (74) that interacts with the Einstein tensor in the Einstein equation. According to (73) the vacuum temperature $T\approx\hbar H$. Therefore, the ratio (74) is rewritten as follows:
Statement 2. If we know the energy density as a function of temperature, then the free energy density is
Checking the Statement 2 is reduced to checking the fact that the function (76) satisfies the following system of equations:
@@ VI Conclussion
was proposed. This equality holds both in the lattice theory of gravity, identical to the one studied here, and in the continuous theory. The relation (80) is justified by the fact that the values under the average transform identically under all symmetry transformations, including discrete PT symmetry. The relation (46), obtained by direct calculation in the continuum limit at zero temperature, is another argument in favor of the validity of the relation (80).
@@
This result is obtained directly in the same way as the result (45) was obtained. The last equality means that in the lattice theory of gravity in the high-temperature phase, space collapses into a point.
However, in the long-wave limit at low temperature, the energy-momentum tensor of the Dirac field is non-zero (46), the quantized field of the tetrad fluctuates weakly and its mean is not equal to zero.
The above means that in the lattice theory of gravity under study there is a phase transition in temperature (possibly more than one). In the high-temperature phase, space is curled up into a point, the average of the energy-momentum tensor is zero, and the discrete PT symmetry is not broken. On the contrary, in the low-temperature phase these quantities are not equal to zero and PT symmetry is broken. The role of the order parameter is played by the average $\langle e^{a}_{\mu}\rangle$, which becomes non-zero in the low-temperature phase. In the low-temperature phase, the process of exponential expansion of space begins, transitioning to a power law of expansion. During the phase transition from the high-temperature to the low-temperature phase, domains with opposite mean values $\langle e^{a}_{\mu}\rangle$ and $\langle\Theta^{a}_{\mu}\rangle$ can be formed. The domain wall separating such domains was studied in Vergeles (2021b)